What is the value today of receiving $2,593.00 per year forever? Assume the first payment is made 8.00 years from today and the discount rate is 6.00%.

Submit

Answer format: Currency: Round to: 2 decimal places.

unanswered

not_submitted

#4

If you are willing to pay $49,200.00 today to receive $4,333.00 per year forever then your required rate of return must be ____%. Assume the first payment is received one year from today.

Submit

Answer format: Percentage Round to: 2 decimal places (Example: 9.24%, % sign required. Will accept decimal format rounded to 4 decimal places (ex: 0.0924))

unanswered

not_submitted

#5

If you are willing to pay $20,149.00 today to receive a perpetuity with the first payment occurring next year then the payment must be $______. Assume a 13.00% discount rate.

Submit

Answer format: Currency: Round to: 2 decimal places.

unanswered

not_submitted

#6

It is _______ for a company to issue equity than debt; it is ________ for an investor to buy equity in a company than debt in the same firm.

a. safer; safer

b. safer; riskier

c. riskier; safer

d. riskier; riskier

e. none of the above

enter a, b, c, d or e
Answer #1

#3

Value of payment 8 year from now:

Present value of annuity of infinite time = cash flow/ interest rate

Cash flow =$2593

Interest rate r= 6%

PV= 2593/0.06

= 43,216.67

Now this present value is for 8 years from now,

So it is a future value for today

N= 8 years

R= 6%

Future Value    = Present Value * (1 + interest rate%)^n

43216.67= PV* (1+0.06)^8

PV= 43216.67/1.5938= $27,115.5